# The theory of implicit operations

### v1

@@ 13. Absolutely closed and primal algebras
The aim of this section is to provide additional criteria to establish whether a pp expansion of a quasivariety is a Beth companion. The two main results are Theorems 13.3 and 13.25 involving absolutely closed algebras and primal algebras, respectively. We will employ these results to determine the Beth companions of torsion-free Abelian groups (13.6), Abelian $\ell$-groups (13.13), MV-algebras (13.17), and varieties of MV-algebras generated by a finite chain (13.26).
@@
We are now ready to illustrate how 13.3 can be applied to describe the Beth companions of concrete classes of algebras.
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{eq}(\mathsf{TFAG})$ such that for all $\bm{A}\in\mathsf{TFAG}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
Let $\varphi_{n}(x,y)=x\thickapprox ny$. Moreover, let $\mathbb{Z}$ and $\mathbb{Q}$ be the additive groups of the integers and the rationals, respectively. The fundamental theorem of finitely generated Abelian groups (see, e.g., [45, Thm. 5.2.3]) implies that every finitely generated torsion-free Abelian group is isomorphic to $\mathbb{Z}^{m}$ for some $m\in\mathbb{N}$. 2.16 implies that $\mathsf{TFAG}$ is generated as a quasivariety by its finitely generated members. Since $\mathbb{Z}\leqslant\mathbb{Q}$, we obtain that $\mathsf{TFAG}$ is also generated as a quasivariety by $\mathbb{Q}$. Observe that for all $\frac{a}{b},\frac{c}{d}\in\mathbb{Q}$ and $n\in\mathbb{Z}^{+}$ we have that $\mathbb{Q}\vDash\varphi_{n}(\frac{a}{b},\frac{c}{d})$ implies $\frac{c}{d}=\frac{a}{nb}$. Therefore, each $\varphi_{n}$ is functional in $\mathbb{Q}$. As $\mathsf{TFAG}$ is the quasivariety generated by $\mathbb{Q}$, 3.11 yields that $\varphi_{n}$ defines a member $f_{n}$ of $\mathsf{imp}_{eq}(\mathsf{TFAG})$. From the definition of $\varphi_{n}$ it follows that the two displays in the statement hold. Moreover, $f_{n}^{\mathbb{Q}}$ is total because every rational can be divided by $n$ in $\mathbb{Q}$. As $\mathsf{TFAG}$ is generated as a quasivariety by $\mathbb{Q}$, 8.11(ii) guarantees that $f_{n}$ is extendable. ∎
@@
Let $f_{2}$ be the member of $\mathsf{ext}_{eq}(\mathsf{TFAG})$ given by 13.7. Since $\mathbb{Z}\leqslant\mathbb{Q}\in\mathsf{TFAG}$ and $f_{2}^{\mathbb{Q}}(1)=\frac{1}{2}\notin\mathbb{Z}$, from 4.10 it follows that $\frac{1}{2}\in\mathsf{d}_{\mathsf{TFAG}}(\mathbb{Z},\mathbb{Q})$-$\mathbb{Z}$. Therefore, $\mathsf{TFAG}$ lacks the strong epimorphism surjectivity property by 6.6. ∎
Let $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}$ be the set of implicit operations given by 13.7. By the same proposition we have $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{TFAG})$. Denote by L{L} the language of groups and let ${L}_{\mathcal{F}}={L}\cup\{d_{n}:n\in\mathbb{Z}^{+}\}$ be an $\mathcal{F}$-expansion of L{L} in which the role of $g_{f_{n}}$ is played by $d_{n}$. Then $\mathsf{DAG}=\mathbb{S}(\mathsf{TFAG}[{L}_{\mathcal{F}}])$ is an equational pp expansion of $\mathsf{TFAG}$. We will prove that it is an equational Beth companion of $\mathsf{TFAG}$. To this end, we rely on the following observation.
@@
Let $\bm{A}\in\mathsf{TFAG}$. As $\mathcal{F}\subseteq\mathsf{ext}_{eq}(\mathsf{TFAG})$, from 10.4 it follows that $\mathsf{DAG}=\mathsf{TFAG}[{L}_{\mathcal{F}}]$. Then $\bm{A}\in\mathsf{DAG}{\upharpoonright}_{{L}}$ if and only if $f_{n}^{\bm{A}}$ is total for every $n\in\mathbb{Z}^{+}$. Therefore, $\bm{A}\in\mathsf{DAG}{\upharpoonright}_{{L}}$ if and only if $\bm{A}$ is divisible. Since $\mathsf{DAG}{\upharpoonright}_{{L}}\subseteq\mathsf{TFAG}$, we conclude that the members of $\mathsf{DAG}{\upharpoonright}_{{L}}$ are exactly the divisible torsion-free Abelian groups. It only remains to prove that every member of $\mathsf{DAG}{\upharpoonright}_{{L}}$ is injective in $\mathsf{TFAG}$. It is well known that the injective members of the variety $\mathsf{AG}$ of Abelian groups are exactly the divisible Abelian groups (see, e.g., [106, Cor. 2.3.2]). Since $\mathsf{TFAG}\subseteq\mathsf{AG}$, from the definition of an injective algebra it follows that every divisible member of $\mathsf{AG}$ that is torsion-free is also injective in $\mathsf{TFAG}$. As all the members of $\mathsf{DAG}{\upharpoonright}_{L}$ are divisible, we conclude they are injective in $\mathsf{TFAG}$. ∎
@@
Let $\Sigma$ be a set of equations axiomatizing the variety of Abelian groups. Since $\varphi_{n}=x\thickapprox ny$ is the equation defining $f_{n}$ (see the proof of 13.7), from 10.4 it follows that $\mathsf{DAG}$ is axiomatized by the set of quasiequations
@@
Lastly, by 13.9 every member of $\mathsf{DAG}{\upharpoonright}_{{L}}$ is injective in $\mathsf{TFAG}$. Then 13.5 implies $\mathsf{DAG}{\upharpoonright}_{{L}}\subseteq\mathsf{TFAG}_{\textsc{ac}}$. Since $\mathsf{DAG}$ is a pp expansion of $\mathsf{TFAG}$, from 13.3 it follows that $\mathsf{DAG}$ is a Beth companion of $\mathsf{TFAG}$ which, moreover, is equational because $\mathsf{DAG}$ is an equational pp expansion of $\mathsf{TFAG}$. ∎
@@
(ii): Assume that $\mathsf{K}$ has the amalgamation property and that $\bm{A}$ is saturated in $\mathsf{K}$. Let $\bm{B}\in\mathsf{K}$ be such that $\bm{A}\leqslant\bm{B}$. To prove that $\bm{A}$ is absolutely closed in $\mathsf{K}$, we need to show that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$. Consider the subalgebra $\bm{D}$ of $\bm{B}$ with universe $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Then $\bm{A}\leqslant\bm{D}\leqslant\bm{B}$. Since $\mathsf{K}$ has the amalgamation property and is a quasivariety by assumption, 4.11 implies $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D})=\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\cap D$. As $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=D$, we obtain $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D})=D$. Furthermore, $\bm{D}\in\mathsf{K}$ because $\bm{D}\leqslant\bm{B}\in\mathsf{K}$ and $\mathsf{K}$ is a quasivariety. It follows that $\bm{A}$ is a $\mathsf{K}$-epic subalgebra of $\bm{D}$. Then the assumption that $\bm{A}$ is saturated in $\mathsf{K}$ and $\bm{D}\in\mathsf{K}$ let us conclude that $\bm{A}=\bm{D}$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$, as desired. ∎
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{eq}({\ell\mathsf{AG}})$ such that for all $\bm{A}\in{\ell\mathsf{AG}}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
Let $\varphi_{n}(x,y)=x\thickapprox ny$. The proof of 13.7 shows that $\varphi_{n}$ defines a unary implicit operation of $\mathsf{TFAG}$. As the class of group reducts of $\ell\mathsf{AG}$} is $\mathsf{TFAG}$ (see [16, Cor. of Thm. XIII.11]), the equation $\varphi_{n}$ defines also a unary $f_{n}\in\mathsf{imp}_{eq}({\ell\mathsf{AG}})$. Clearly, the two displays in the statement hold for $f_{n}$. Therefore, it only remains to show that $f_{n}$ is extendable. Recall from [15, Lem. 1, p. 317] that the members of ${\ell\mathsf{AG}}_{\textsc{rfsi}}$ are nontrivial and linearly ordered. Moreover, the class of nontrivial linearly ordered members of ${\ell\mathsf{AG}}_{\textsc{rfsi}}$ is $\mathbb{U}(\mathbb{Q})$, where $\mathbb{Q}$ denotes the additive group of the rationals equipped with the lattice structure induced by the standard order of $\mathbb{Q}$ by [60]. Therefore, the Subdirect Decomposition Theorem 2.9 yields that $\ell\mathsf{AG}$} is generated as a quasivariety by $\mathbb{Q}$. Arguing as in the proof of 13.7, we conclude that each $f_{n}$ is extendable. ∎
@@
Analogous to the proof of 13.8 with the sole difference that $\mathbb{Q}$ is the algebra employed in the proof of 13.14 and $\mathbb{Z}$ the subalgebra of $\mathbb{Q}$ whose universe is the set of integers. ∎
Set $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}$ be the set of implicit operations given by 13.14. By the same proposition we have $\mathcal{F}\subseteq\mathsf{ext}_{pp}(\mathsf{TFAG})$. Let ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{{\ell\mathsf{AG}}}$. Then ${\ell\mathsf{DAG}}=\mathbb{S}({\ell\mathsf{AG}}[{L}_{\mathcal{F}}])$ is an equational pp expansion of $\ell\mathsf{AG}$}.
@@
Since $\mathcal{F}\subseteq\mathsf{ext}_{eq}({\ell\mathsf{AG}})$ and $\ell\mathsf{AG}$} is a variety, 10.3(iii) implies that $\ell\mathsf{DAG}$} is also a variety. Moreover, from 10.4 it follows that ${\ell\mathsf{DAG}}={\ell\mathsf{AG}}[{L}_{\mathcal{F}}]$. Therefore, ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}$ is the class of divisible Abelian $\ell$-groups. Then [2, Thm. 2.1] yields that ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}$ is the class of members of $\ell\mathsf{AG}$} that are saturated in $\ell\mathsf{AG}$}. Since $\ell\mathsf{Ab}$ has the amalgamation property (see [98, Thm. 2.3]), we can apply 13.12(ii), obtaining ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}\subseteq{\ell\mathsf{AG}}_{\textsc{ac}}$. Thus, 13.3 implies that $\ell\mathsf{DAG}$} is a Beth companion of $\ell\mathsf{AG}$} which, moreover, is equational because $\ell\mathsf{DAG}$} is an equational pp expansion of $\ell\mathsf{AG}$}. ∎
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{eq}(\mathsf{MV})$ such that for all $\bm{A}\in\mathsf{MV}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
From 13.19(iii) it follows that $\varphi_{n}$ is functional and total in [0,1]{[0,1]}. Since $\mathsf{MV}=\mathbb{Q}({[0,1]})$ (see, e.g., [56, p. 84]), 3.11 and 8.11(ii) imply that $\varphi_{n}$ defines a unary $f_{n}\in\mathsf{ext}_{eq}(\mathsf{MV})$. Lastly, as $\varphi_{n}$ defines $f_{n}$, the two displays in the statement hold. ∎
@@
Let $f_{2}$ be the member of $\mathsf{ext}_{eq}(\mathsf{MV})$ given by 13.18. Moreover, let $\bm{A}$ be the subalgebra of [0,1]{[0,1]} with universe $\{0,1\}$. Since $\bm{A}\leqslant{[0,1]}\in\mathsf{MV}$ and $f_{2}^{[0,1]}(1)=\frac{1}{2}\notin A$, from 4.10 it follows that $\frac{1}{2}\in\mathsf{d}_{\mathsf{MV}}(\bm{A},{[0,1]})-A$. Therefore, $\mathsf{MV}$ lacks the strong epimorphism surjectivity property by 6.6. ∎
@@
Let $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}\subseteq\mathsf{ext}_{eq}(\mathsf{MV})$ be the family of operations given by 13.18. Moreover, let $d_{n}$ be a unary function symbol for each $n\in\mathbb{Z}^{+}$. Then the language ${L}_{\mathcal{F}}={L}_{\mathsf{MV}}\cup\{d_{n}:n\in\mathbb{Z}^{+}\}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{MV}}$ in which the role of $g_{f_{n}}$ is played by $d_{n}$. From 10.4 and the fact that each $f_{n}$ is defined by the conjunction of equations $\varphi_{n}$ in the proof of 13.18 it follows that $\mathbb{S}(\mathsf{MV}[{L}_{\mathcal{F}}])=\mathsf{MV}[{L}_{\mathcal{F}}]$ is an equational pp expansion of $\mathsf{MV}$ axiomatized by the axioms of MV-algebras plus the equations $n.d_{n}(x)\thickapprox x$ and $d_{n}(x)\odot((n-1).d_{n}(x))\thickapprox 0$ for $n\in\mathbb{Z}^{+}$. Clearly, every member of $\mathsf{MV}[{L}_{\mathcal{F}}]$ is a DMV-algebra. On the other hand, every DMV-algebra can be obtained by adding the implicit operations $f_{n}$ to its MV-algebra reduct, which belongs to $\mathsf{MV}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{MV}}}$. Therefore, $\mathsf{MV}[{L}_{\mathcal{F}}]$ coincides with the variety $\mathsf{DMV}$ of DMV-algebras.
In view of 13.3, to show that $\mathsf{DMV}$ is a Beth companion of $\mathsf{MV}$, it suffices to prove that $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}\subseteq\mathsf{MV}_{\textsc{ac}}$. The definition of $\mathsf{DMV}$ yields that the members of $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}$ are divisible MV-algebras. Every divisible MV-algebra is saturated in $\mathsf{MV}$ (see [47, Thm. 3.18(ii)]) and $\mathsf{MV}$ has the amalgamation property (see [96, p. 91]). Therefore, from 13.12(ii) it follows that $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}\subseteq\mathsf{MV}_{\textsc{ac}}$. Then 13.3 yields that $\mathsf{DMV}$ is a Beth companion of $\mathsf{MV}$ which, moreover, is equational because $\mathsf{DMV}$ is an equational pp expansion of $\mathsf{MV}$. ∎
@@
The structure theory of varieties generated by a primal algebra is very rich, as a consequence of the fact that these are precisely the varieties categorically equivalent to the variety of Boolean algebras [70] (see also [41]). In particular, since the variety of Boolean algebras has the surjective epimorphism property (see 7.5) and this property is preserved by categorical equivalences between varieties by 6.4, we deduce the following.

### v2

@@ 13. Absolutely closed and primal algebras
The aim of this section is to provide additional criteria to establish whether a pp expansion of a quasivariety is a Beth companion. The two main results are Theorems 13.3 and 13.25 involving absolutely closed algebras and primal algebras, respectively. We will employ these results to determine the Beth companions of torsion-free Abelian groups (Example 13.6), Abelian $\ell$-groups (Example 13.13), MV-algebras (Example 13.17), and varieties of MV-algebras generated by a finite chain (Example 13.26).
@@
We are now ready to illustrate how Theorem 13.3 can be applied to describe the Beth companions of concrete classes of algebras.
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{\textsc{eq}}(\mathsf{TFAG})$ such that for all $\bm{A}\in\mathsf{TFAG}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
Let $\varphi_{n}(x,y)=x\thickapprox ny$. Moreover, let $\mathbb{Z}$ and $\mathbb{Q}$ be the additive groups of the integers and the rationals, respectively. The fundamental theorem of finitely generated Abelian groups (see, e.g., [45, Thm. 5.2.3]) implies that every finitely generated torsion-free Abelian group is isomorphic to $\mathbb{Z}^{m}$ for some $m\in\mathbb{N}$. Proposition 2.16 implies that $\mathsf{TFAG}$ is generated as a quasivariety by its finitely generated members. Since $\mathbb{Z}\leqslant\mathbb{Q}$, we obtain that $\mathsf{TFAG}$ is also generated as a quasivariety by $\mathbb{Q}$. Observe that for all $\frac{a}{b},\frac{c}{d}\in\mathbb{Q}$ and $n\in\mathbb{Z}^{+}$ we have that $\mathbb{Q}\vDash\varphi_{n}(\frac{a}{b},\frac{c}{d})$ implies $\frac{c}{d}=\frac{a}{nb}$. Therefore, each $\varphi_{n}$ is functional in $\mathbb{Q}$. As $\mathsf{TFAG}$ is the quasivariety generated by $\mathbb{Q}$, Corollary 3.11 yields that $\varphi_{n}$ defines a member $f_{n}$ of $\mathsf{imp}_{\textsc{eq}}(\mathsf{TFAG})$. From the definition of $\varphi_{n}$ it follows that the two displays in the statement hold. Moreover, $f_{n}^{\mathbb{Q}}$ is total because every rational can be divided by $n$ in $\mathbb{Q}$. As $\mathsf{TFAG}$ is generated as a quasivariety by $\mathbb{Q}$, Proposition 8.11(ii) guarantees that $f_{n}$ is extendable. ∎
@@
Let $f_{2}$ be the member of $\mathsf{ext}_{\textsc{eq}}(\mathsf{TFAG})$ given by Proposition 13.7. Since $\mathbb{Z}\leqslant\mathbb{Q}\in\mathsf{TFAG}$ and $f_{2}^{\mathbb{Q}}(1)=\frac{1}{2}\notin\mathbb{Z}$, from Theorem 4.10 it follows that $\frac{1}{2}\in\mathsf{d}_{\mathsf{TFAG}}(\mathbb{Z},\mathbb{Q})$-$\mathbb{Z}$. Therefore, $\mathsf{TFAG}$ lacks the strong epimorphism surjectivity property by Proposition 6.6. ∎
Let $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}$ be the set of implicit operations given by Proposition 13.7. By the same proposition we have $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{TFAG})$. Denote by L{L} the language of groups and let ${L}_{\mathcal{F}}={L}\cup\{d_{n}:n\in\mathbb{Z}^{+}\}$ be an $\mathcal{F}$-expansion of L{L} in which the role of $g_{f_{n}}$ is played by $d_{n}$. Then $\mathsf{DAG}=\mathbb{S}(\mathsf{TFAG}[{L}_{\mathcal{F}}])$ is an equational pp expansion of $\mathsf{TFAG}$. We will prove that it is an equational Beth companion of $\mathsf{TFAG}$. To this end, we rely on the following observation.
@@
Let $\bm{A}\in\mathsf{TFAG}$. As $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{TFAG})$, from Theorem 10.4 it follows that $\mathsf{DAG}=\mathsf{TFAG}[{L}_{\mathcal{F}}]$. Then $\bm{A}\in\mathsf{DAG}{\upharpoonright}_{{L}}$ if and only if $f_{n}^{\bm{A}}$ is total for every $n\in\mathbb{Z}^{+}$. Therefore, $\bm{A}\in\mathsf{DAG}{\upharpoonright}_{{L}}$ if and only if $\bm{A}$ is divisible. Since $\mathsf{DAG}{\upharpoonright}_{{L}}\subseteq\mathsf{TFAG}$, we conclude that the members of $\mathsf{DAG}{\upharpoonright}_{{L}}$ are exactly the divisible torsion-free Abelian groups. It only remains to prove that every member of $\mathsf{DAG}{\upharpoonright}_{{L}}$ is injective in $\mathsf{TFAG}$. It is well known that the injective members of the variety $\mathsf{AG}$ of Abelian groups are exactly the divisible Abelian groups (see, e.g., [106, Cor. 2.3.2]). Since $\mathsf{TFAG}\subseteq\mathsf{AG}$, from the definition of an injective algebra it follows that every divisible member of $\mathsf{AG}$ that is torsion-free is also injective in $\mathsf{TFAG}$. As all the members of $\mathsf{DAG}{\upharpoonright}_{L}$ are divisible, we conclude they are injective in $\mathsf{TFAG}$. ∎
@@
Let $\Sigma$ be a set of equations axiomatizing the variety of Abelian groups. Since $\varphi_{n}=x\thickapprox ny$ is the equation defining $f_{n}$ (see the proof of Proposition 13.7), from Theorem 10.4 it follows that $\mathsf{DAG}$ is axiomatized by the set of quasiequations
@@
Lastly, by Proposition 13.9 every member of $\mathsf{DAG}{\upharpoonright}_{{L}}$ is injective in $\mathsf{TFAG}$. Then Proposition 13.5 implies $\mathsf{DAG}{\upharpoonright}_{{L}}\subseteq\mathsf{TFAG}_{\textsc{ac}}$. Since $\mathsf{DAG}$ is a pp expansion of $\mathsf{TFAG}$, from Theorem 13.3 it follows that $\mathsf{DAG}$ is a Beth companion of $\mathsf{TFAG}$ which, moreover, is equational because $\mathsf{DAG}$ is an equational pp expansion of $\mathsf{TFAG}$. ∎
@@
(ii): Assume that $\mathsf{K}$ has the amalgamation property and that $\bm{A}$ is saturated in $\mathsf{K}$. Let $\bm{B}\in\mathsf{K}$ be such that $\bm{A}\leqslant\bm{B}$. To prove that $\bm{A}$ is absolutely closed in $\mathsf{K}$, we need to show that $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$. Consider the subalgebra $\bm{D}$ of $\bm{B}$ with universe $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})$. Then $\bm{A}\leqslant\bm{D}\leqslant\bm{B}$. Since $\mathsf{K}$ has the amalgamation property and is a quasivariety by assumption, Proposition 4.11 implies $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D})=\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})\cap D$. As $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=D$, we obtain $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{D})=D$. Furthermore, $\bm{D}\in\mathsf{K}$ because $\bm{D}\leqslant\bm{B}\in\mathsf{K}$ and $\mathsf{K}$ is a quasivariety. It follows that $\bm{A}$ is a $\mathsf{K}$-epic subalgebra of $\bm{D}$. Then the assumption that $\bm{A}$ is saturated in $\mathsf{K}$ and $\bm{D}\in\mathsf{K}$ let us conclude that $\bm{A}=\bm{D}$. Therefore, $\mathsf{d}_{\mathsf{K}}(\bm{A},\bm{B})=A$, as desired. ∎
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{\textsc{eq}}({\ell\mathsf{AG}})$ such that for all $\bm{A}\in{\ell\mathsf{AG}}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
Let $\varphi_{n}(x,y)=x\thickapprox ny$. The proof of Proposition 13.7 shows that $\varphi_{n}$ defines a unary implicit operation of $\mathsf{TFAG}$. As the class of group reducts of $\ell\mathsf{AG}$} is $\mathsf{TFAG}$ (see [16, Cor. of Thm. XIII.11]), the equation $\varphi_{n}$ defines also a unary $f_{n}\in\mathsf{imp}_{\textsc{eq}}({\ell\mathsf{AG}})$. Clearly, the two displays in the statement hold for $f_{n}$. Therefore, it only remains to show that $f_{n}$ is extendable. Recall from [15, Lem. 1, p. 317] that the members of ${\ell\mathsf{AG}}_{\textsc{rfsi}}$ are nontrivial and linearly ordered. Moreover, the class of nontrivial linearly ordered members of ${\ell\mathsf{AG}}_{\textsc{rfsi}}$ is $\mathbb{U}(\mathbb{Q})$, where $\mathbb{Q}$ denotes the additive group of the rationals equipped with the lattice structure induced by the standard order of $\mathbb{Q}$ by [60]. Therefore, the Subdirect Decomposition Theorem 2.9 yields that $\ell\mathsf{AG}$} is generated as a quasivariety by $\mathbb{Q}$. Arguing as in the proof of Proposition 13.7, we conclude that each $f_{n}$ is extendable. ∎
@@
Analogous to the proof of Corollary 13.8 with the sole difference that $\mathbb{Q}$ is the algebra employed in the proof of Proposition 13.14 and $\mathbb{Z}$ the subalgebra of $\mathbb{Q}$ whose universe is the set of integers. ∎
Set $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}$ be the set of implicit operations given by Proposition 13.14. By the same proposition we have $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{pp}}(\mathsf{TFAG})$. Let ${L}_{\mathcal{F}}$ be an $\mathcal{F}$-expansion of ${L}_{{\ell\mathsf{AG}}}$. Then ${\ell\mathsf{DAG}}=\mathbb{S}({\ell\mathsf{AG}}[{L}_{\mathcal{F}}])$ is an equational pp expansion of $\ell\mathsf{AG}$}.
@@
Since $\mathcal{F}\subseteq\mathsf{ext}_{\textsc{eq}}({\ell\mathsf{AG}})$ and $\ell\mathsf{AG}$} is a variety, Theorem 10.3(iii) implies that $\ell\mathsf{DAG}$} is also a variety. Moreover, from Theorem 10.4 it follows that ${\ell\mathsf{DAG}}={\ell\mathsf{AG}}[{L}_{\mathcal{F}}]$. Therefore, ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}$ is the class of divisible Abelian $\ell$-groups. Then [2, Thm. 2.1] yields that ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}$ is the class of members of $\ell\mathsf{AG}$} that are saturated in $\ell\mathsf{AG}$}. Since $\ell\mathsf{Ab}$ has the amalgamation property (see [98, Thm. 2.3]), we can apply Proposition 13.12(ii), obtaining ${\ell\mathsf{DAG}}{\upharpoonright}_{{L}_{\ell\mathsf{AG}}}\subseteq{\ell\mathsf{AG}}_{\textsc{ac}}$. Thus, Theorem 13.3 implies that $\ell\mathsf{DAG}$} is a Beth companion of $\ell\mathsf{AG}$} which, moreover, is equational because $\ell\mathsf{DAG}$} is an equational pp expansion of $\ell\mathsf{AG}$}. ∎
@@
For each $n\in\mathbb{Z}^{+}$ there exists a unary $f_{n}\in\mathsf{ext}_{\textsc{eq}}(\mathsf{MV})$ such that for all $\bm{A}\in\mathsf{MV}$ and $a\in\mathsf{dom}(f_{n}^{\bm{A}})$,
@@
From Claim 13.19(iii) it follows that $\varphi_{n}$ is functional and total in [0,1]{[0,1]}. Since $\mathsf{MV}=\mathbb{Q}({[0,1]})$ (see, e.g., [56, p. 84]), Corollary 3.11 and Proposition 8.11(ii) imply that $\varphi_{n}$ defines a unary $f_{n}\in\mathsf{ext}_{\textsc{eq}}(\mathsf{MV})$. Lastly, as $\varphi_{n}$ defines $f_{n}$, the two displays in the statement hold. ∎
@@
Let $f_{2}$ be the member of $\mathsf{ext}_{\textsc{eq}}(\mathsf{MV})$ given by Proposition 13.18. Moreover, let $\bm{A}$ be the subalgebra of [0,1]{[0,1]} with universe $\{0,1\}$. Since $\bm{A}\leqslant{[0,1]}\in\mathsf{MV}$ and $f_{2}^{[0,1]}(1)=\frac{1}{2}\notin A$, from Theorem 4.10 it follows that $\frac{1}{2}\in\mathsf{d}_{\mathsf{MV}}(\bm{A},{[0,1]})-A$. Therefore, $\mathsf{MV}$ lacks the strong epimorphism surjectivity property by Proposition 6.6. ∎
@@
Let $\mathcal{F}=\{f_{n}:n\in\mathbb{Z}^{+}\}\subseteq\mathsf{ext}_{\textsc{eq}}(\mathsf{MV})$ be the family of operations given by Proposition 13.18. Moreover, let $d_{n}$ be a unary function symbol for each $n\in\mathbb{Z}^{+}$. Then the language ${L}_{\mathcal{F}}={L}_{\mathsf{MV}}\cup\{d_{n}:n\in\mathbb{Z}^{+}\}$ is an $\mathcal{F}$-expansion of ${L}_{\mathsf{MV}}$ in which the role of $g_{f_{n}}$ is played by $d_{n}$. From Theorem 10.4 and the fact that each $f_{n}$ is defined by the conjunction of equations $\varphi_{n}$ in the proof of Proposition 13.18 it follows that $\mathbb{S}(\mathsf{MV}[{L}_{\mathcal{F}}])=\mathsf{MV}[{L}_{\mathcal{F}}]$ is an equational pp expansion of $\mathsf{MV}$ axiomatized by the axioms of MV-algebras plus the equations $n.d_{n}(x)\thickapprox x$ and $d_{n}(x)\odot((n-1).d_{n}(x))\thickapprox 0$ for $n\in\mathbb{Z}^{+}$. Clearly, every member of $\mathsf{MV}[{L}_{\mathcal{F}}]$ is a DMV-algebra. On the other hand, every DMV-algebra can be obtained by adding the implicit operations $f_{n}$ to its MV-algebra reduct, which belongs to $\mathsf{MV}[{L}_{\mathcal{F}}]{\upharpoonright}_{{L}_{\mathsf{MV}}}$. Therefore, $\mathsf{MV}[{L}_{\mathcal{F}}]$ coincides with the variety $\mathsf{DMV}$ of DMV-algebras.
In view of Theorem 13.3, to show that $\mathsf{DMV}$ is a Beth companion of $\mathsf{MV}$, it suffices to prove that $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}\subseteq\mathsf{MV}_{\textsc{ac}}$. The definition of $\mathsf{DMV}$ yields that the members of $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}$ are divisible MV-algebras. Every divisible MV-algebra is saturated in $\mathsf{MV}$ (see [47, Thm. 3.18(ii)]) and $\mathsf{MV}$ has the amalgamation property (see [96, p. 91]). Therefore, from Proposition 13.12(ii) it follows that $\mathsf{DMV}{\upharpoonright}_{\mathsf{MV}}\subseteq\mathsf{MV}_{\textsc{ac}}$. Then Theorem 13.3 yields that $\mathsf{DMV}$ is a Beth companion of $\mathsf{MV}$ which, moreover, is equational because $\mathsf{DMV}$ is an equational pp expansion of $\mathsf{MV}$. ∎
@@
The structure theory of varieties generated by a primal algebra is very rich, as a consequence of the fact that these are precisely the varieties categorically equivalent to the variety of Boolean algebras [70] (see also [41]). In particular, since the variety of Boolean algebras has the surjective epimorphism property (see Theorem 7.5) and this property is preserved by categorical equivalences between varieties by Remark 6.4, we deduce the following.